OpenStudy (anonymous):

helppp .Medal & Fan. Simplify the complex fraction. 4/(x+3)/(1/x+3) A. 12x+4/x^2+3x B. 4x/3x+9 C. 4x/3x^2+10x+3 D. x^2+3x/12x+4

OpenStudy (anonymous):

@Abdulhameed could you helpp (:

OpenStudy (anonymous):

can u rewrite it plz

OpenStudy (jdoe0001):

\(\bf \cfrac{\frac{4}{x+3}}{\frac{1}{x+3}}\qquad recall\implies \cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\cdot \cfrac{d}{c}\qquad thus\\ \quad \\ \cfrac{\frac{4}{x+3}}{\frac{1}{x+3}}\implies \cfrac{4}{\cancel{x+3}}\cdot \cfrac{\cancel{x+3}}{1}\)

OpenStudy (anonymous):

 ??